# Measuring Casimir Force Across a Superconducting Transition

### v1

@@ Abstract
The Casimir effect and superconductivity are foundational quantum phenomena whose interaction remains an open question in physics. How Casimir forces behave across a superconducting transition remains unresolved, owing to the experimental difficulty of achieving alignment, cryogenic environments, and isolating small changes from competing effects. This question carries implications for electron physics, quantum gravity, and high-temperature superconductivity. Here we demonstrate an on-chip superconducting platform that overcomes these challenges, achieving one of the most parallel Casimir configurations to date. Our microchip-based cavities achieve unprecedented area-to-separation ratio between plates, exceeding previous Casimir experiments by orders of magnitude and generating the strongest Casimir forces yet between compliant surfaces. Scanning tunneling microscopy (STM) is used for the first time to directly detect the resonant motion of a suspended membrane, with subatomic precision in both lateral positioning and displacement. Such precision measurements across a superconducting transition allow for the suppression of all van der Waals, electrostatic, and thermal effects. Preliminary measurements suggest superconductivity-dependent shifts in the Casimir force, motivating further investigation and comparison with theories. By uniting extreme parallelism, nanomechanics, and STM readout, our platform opens a new experimental frontier at the intersection of Casimir physics and superconductivity.
@@ I Interplay of Casimir Forces and Superconductivity
Figure 1: Superconducting Casimir Force (a) Schematic of a superconducting Casimir cavity, where two plates of area $A$ are separated by a gap d. Inset shows a superconducting transition of the plates. (b) Key challenges in measuring the Casimir effect across $T_{C}$, including achieving high parallelism using compliant plates, and ensuring minimally invasive measurement across $T_{C}$. Additionally, isolating the Casimir force from electrostatic and measurement-induced contributions remains a significant challenge. (c) The Casimir-superconductivity regime links to long-standing controversy in the low-frequency response optical response of metals (Drude vs. plasma), non-classical gravity, and high-$T_{C}$ superconductivity in layered cuprates.
The Casimir effect and superconductivity are two landmark quantum phenomena, each showcasing striking macroscopic behavior rooted in quantum mechanics. The Casimir effect [1], predicted by Hendrik Casimir in 1948, arises from vacuum fluctuations of electromagnetic fields and manifests as an attractive force between two closely spaced, uncharged surfaces. In Fig. 1a, we show Casimir’s original theoretical setup: two perfectly conducting, parallel plates experience forces that scale as $F\propto A/d^{4}$, where $A$ is area and $d$ is separation. These extreme geometries amplify quantum fluctuations into measurable macroscopic forces but require precise alignment and sensitive detection to isolate Casimir effects from competing interactions. Although these challenges delayed experimental confirmation for nearly half a century [2], Casimir forces are now known to dominate nanoscale systems as separations shrink.
Superconductivity, discovered in 1911, occurs when electrons form Cooper pairs [3] below a critical temperature ($T_{C}$), allowing them to move without electrical resistance (Fig. 1a inset). Above $T_{C}$, electrons act as individual particles; below it, they pair into macroscopic quantum states with coherence lengths spanning hundreds of nanometers. This sharp transition into the mesoscopic regime makes superconductors one of the most striking phenomena in condensed matter physics.
Although Casimir’s original theory assumed perfectly conducting plates, an intriguing question remains: how do Casimir forces behave when the plates become superconducting? Despite the fundamental importance of both phenomena, their interplay remains experimentally elusive. The sudden onset of superconductivity at $T_{C}$ offers an opportunity to toggle this interaction on and off, enabling precise differential measurements [4, 5]. Realizing this requires overcoming three primary challenges, outlined in Fig. 1b.
1. Achieving nanometer-scale parallelism at cryogenic temperatures demands exceptional mechanical precision and active feedback control [6, 7], as even slight misalignments distort the Casimir force.
@@
2. Measurements above and below $T_{C}$ must be minimally invasive (avoiding degradation of the Casimir signal and the superconducting state).
@@
3. Casimir forces must be isolated from competing effects; including thermal expansion, material changes at $T_{C}$ [8, 9, 10], electrostatics [11, 12, 13], and artifacts from the measurement itself.
These challenges — exacerbated in cryogenic setups with limited space and complex infrastructure — have kept this regime largely unexplored, leaving the Casimir-superconductivity intersection an open frontier in physics. In this work, we overcome these barriers using a microchip-based platform that achieves ultra-parallel plate configurations and generates record-high Casimir forces. Our platform features free-standing superconducting membranes, whose dynamics across $T_{C}$ are probed using scanning tunneling microscopy (STM). STM’s atomic-scale precision enables minimally invasive, highly localized readout, allowing systematic suppression of van der Waals, electrostatic, and thermal effects. This provides an exceptionally sensitive detection scheme, offering a first step toward uncovering how superconductivity influences the Casimir force and uniting these two macroscopic quantum phenomena.
@@
The following sections present the theoretical context of Casimir–superconductivity interactions, describe the fabrication of our on-chip cavities, and introduce the STM-based detection method. We then outline protocols for eliminating non-Casimir contributions and report calibrated frequency sweeps across $T_{C}$ to probe superconductivity-induced shifts in the Casimir force.
II Theoretical Implications and Open Questions
Understanding how superconductivity influences the Casimir effect could address foundational questions in condensed matter physics, high-temperature superconductivity, and perhaps non-classical gravity (Fig. 1c). The most established framework, the Drude vs. plasma model debate [14, 15], examines how a metal’s optical properties affect Casimir forces. This debate, unresolved for over two decades, centers on whether dissipation via electron scattering (Drude) or dissipationless electron gas (plasma) better predicts metallic Casimir forces. Observing how the Casimir force changes across $T_{C}$ could provide valuable insights into resolving this puzzle, with higher TC enhancing the effects. When a metal transitions to a superconductor, its optical response changes at frequencies on the order of $k_{B}T_{C}/\hbar$. However, the Casimir interaction spans much higher frequencies, up to $\omega_{c}=c/d$ for nanometer-scale separations, suggesting a small impact from superconductivity. Such experiments were attempted twice with no changes in Casimir force across $T_{C}$ [16, 17] due to insufficient sensitivity to either model [14].
Some speculative theories propose superconductors in Casimir configurations may interact with gravity in ways that depart from classical general relativity [18, 19, 20, 21, 22]. While intriguing, these proposals remain unverified and require further theoretical and experimental scrutiny via superconducting Casimir experiments.
Another open question is whether Casimir forces influence $T_{C}$. Prior studies detected no shift in $T_{C}$ within 12 $\mu$K for lead (Pb) plates separated by a 70nm gap in a plate–plate configuration[23]. High-$T_{C}$ superconductors, however, present a unique case. Their layered structures form intrinsic Casimir cavities, where the associated energy may generate internal pressures exceeding 100atm — comparable to their superconducting condensation energies. Recent theories suggest that Casimir interactions in multilayered cuprates may help stabilize high-$T_{C}$ superconductivity, potentially addressing one of the most enduring and impactful open problems in physics.[24, 25]. Notably, 1960s experiments showed that alternating superconducting and insulating layers enhanced $T_{C}$, hinting at a related mechanism [26, 27].
Whether resolving the Drude-plasma debate, probing high-$T_{C}$ mechanisms, or testing faint gravitational effects, the interplay between superconductivity and Casimir forces offers profound opportunities to advance fundamental physics and materials science with table-top experiments. Yet, experimental challenges have kept this frontier of quantum physics largely unexplored.
III Ultra-Parallel Superconducting Casimir Cavities
Over the past decade, a variety of on-chip Casimir geometries have been explored [28, 29, 30, 16, 31, 23, 32, 33]. Most conventional setups use a sphere–plate geometry for easier alignment between two surfaces. While plate–plate configurations more closely follow Casimir’s original theory, they require stringent parallelism and mechanical stability, making them rare [6, 7, 29, 23]. To date, no superconducting Casimir experiments have reported measurable force shifts across $T_{C}$ [16, 17, 34, 23].
@@ III Ultra-Parallel Superconducting Casimir Cavities
Figure 2: Achieving extreme parallelism on-chip. (a) Colorized SEM of a microchip with suspended NbTiN (square) membranes above a NbTiN backgate; inset shows the layered structure. Membranes feature micron-scale holes for selective plasma etching of the a-Si layer. A $1~{}\mu\rm m$ SiO2 layer electrically isolates the Casimir cavity. (b) A central hole in the top membrane offers visual access to the gap. (c) Close-up of 190nm gaps sustained over millimeter-scale areas, with $\sim$130pm sag kept taut by high-stress NbTiN. (d) Comparison of area-over-gap ratios for published Casimir setups. Here, the ideal Casimir force between perfect conductors is used as a consistent metric to quantify effective parallelism. Blue squares: plate-plate ($F\propto A/d^{4}$); red circles: sphere-plate ($F\propto R/d^{3}$). Our chip platform’s parallelism surpasses prior cryogenic and room-temperature experiments, achieving the highest Casimir force between compliant surfaces to date. See Supplementary Information J for data sources.
Maximizing sensitivity to superconductivity-induced Casimir shifts requires both large Casimir pressures $F_{\text{Cas}}\propto A/d^{4}$ and high mechanical compliance. Thinner, more flexible membranes respond more sensitively to force variations but are also more prone to stiction and misalignment, making extreme parallelism harder to maintain. This trade-off defines a key constraint: optimize parallelism for large Casimir forces while maintaining compliance for detectable frequency shifts.
We overcome this trade-off by nanofabricating Casimir cavities directly on silicon chips using high-stress superconducting films. Our monolithic plate–plate design suspends a $709\times 709~{}\mu\rm m^{2}$ niobium titanium nitride (NbTiN) membrane above a fixed NbTiN backplate with gaps as small as 190 nm (Fig. 2). High tensile stress ($\approx$ 680 MPa) keeps the membrane taut and stable, preventing collapse from stiction or adhesion. The cavity is formed by depositing 155 nm-thick NbTiN layers above and below a sacrificial amorphous silicon (a-Si) layer, which is delicately removed via a highly selective SF6 plasma etch that preserves the NbTiN’s 1–2 nm surface roughness (see Supplementary Information D). After suspension, an adiabatic return to lab conditions from cryogenic vacuum environments prevents collapse, enabling high-yield fabrication. Gap size is tunable by adjusting a-Si thickness. Successfully suspended membranes exhibit fundamental frequencies near 253 kHz in room temperature, confirmed via laser Doppler vibrometry, and near 350 kHz in cryogenic temperature, confirmed by STM measurement.
@@
The 190 nm gap, smallest sustained over millimeter-scale areas, is a large enough separation to avoid complex van der Waals interactions and roughness-induced noise, both of which become significant below 100 nm [35]. Our membranes combine high tensile stress, a superconducting transition temperature ($T_{C}\approx 14.2$ K), and atomically smooth surfaces ($\approx$ 1 nm rms roughness), all crucial for precise Casimir measurements. Achieving both high $T_{C}$ and high tensile stress is nontrivial, as stress typically suppresses superconductivity. Through material optimization, we preserve mechanical performance while ensuring a $T_{C}$ close to bulk NbTiN. This is relevant for superconducting Casimir theories, where higher $T_{C}$ improves spectral overlap between the superconducting gap and Casimir-relevant modes, enhancing the chance of measurable effects.
This fixed-gap design eliminates the need for active alignment, maintaining near-constant separation during temperature sweeps. The main deviation arises from a $\sim$142 picometer sag over a 709 micron span — caused primarily by Casimir attraction — yielding an angular misalignment of just $\sim$2×$\times 10^{-5}$ degrees. While active alignment of two surfaces via precision stages offers tunability, our chip-based platform achieves far superior parallelism and passive stability; ideal for temperature-based relative measurements. As a control, we fabricate an identical device with a 1.25 $\mu$m gap, where Casimir forces are negligible (see Supplementary Information H). This allows us to disentangle Casimir-induced shifts from those caused by thermal expansion, material changes across $T_{C}$, or changes in STM readout.
@@
Fig. 2 illustrates the extreme aspect-ratio parallelism achieved on-chip. A colorized SEM image (Fig. 2a) shows suspended NbTiN membranes with micron-scale release holes. The inset shows a side-view schematic of the Casimir cavity on a silicon substrate. A central hole (Fig. 2b) offers visual confirmation of suspension, though it is omitted in actual experiments where vibrational modes are confirmed via laser Doppler vibrometry. Close-ups (Fig. 2c) reveal 190 nm gaps maintained over millimeter-scale areas. Fig. 2d highlights our platform’s ability to achieve extreme parallelism, defined by an unprecedented area-to-separation ratio compared to previous published Casimir experiments, using the ideal Casimir force as a metric of effective parallelism. Our system’s parallelism surpasses all previous published Casimir experiments (room-temperature and cryogenic) by orders of magnitude. Beyond small separations, our high-stress NbTiN membranes form high-Q mechanical resonators, enabling sensitive detection of Casimir shifts.
@@
Materials engineering is central to this work. High tensile stress in NbTiN membranes prevents collapse while enabling high-Q mechanical modes for precision force sensing via dissipation dilution [36]. Years of optimization allowed us to achieve both high $T_{C}$ and resilience to fabrication chemicals (see Supplementary Information B). The NbTiN stress also enhances membrane vibrational stability, making it possible to interaction with atomic-scale probes like STM. Crucially, NbTiN’s minimal oxidation allows effective in-situ annealing (to allow STM tunneling), unlike aluminum, which forms thick ($\sim$ 2-5 nm) oxides that are difficult to remove. These properties make our platform not only well-suited for generating strong Casimir forces but also for high-precision STM interrogation, as detailed in the next section.
IV Experimental Setup and Measurement Protocol
@@
We use a scanning tunneling microscope (STM) to measure the mechanical resonance of suspended NbTiN membranes across the superconducting transition temperature. The STM tip is positioned with sub-nanometer precision, enabling precision detection of membrane motion. By approaching within 0.3–1.0 nm of the surface, the tip tracks membrane oscillations of 30–100 pm. The tunneling current drops by a factor of 10 for every 0.1 nm increase in separation — changing by 1000$\times$ over a 0.3 nm atomic diameter — making STM exquisitely sensitive to atomic-scale motion. During measurements, the tip remains confined within a $\sim$10 pm lateral region — 30$\times$ smaller than a Nb atom — on a $709\times 709~{}\mu$m2 membrane, spanning 12 orders of magnitude in scale. This localization enables precise cancellations of the tip’s influence on the membrane via van der Waal and electrostatic forces. Its minimal interaction area reduces disruption to the membrane’s superconductivity and enables direct measurement of the superconducting gap.
@@
Fig. 3a shows the experimental setup. Measurements are conducted in ultra-high vacuum ($\sim 10^{-10}$ mbar) and cryogenic temperatures (4.45–14.45 K). A suspended NbTiN membrane vibrates above a fixed NbTiN backgate, driven by an AC+DC voltage (AC for actuation, DC to cancel electrostatics). Scanning tunneling microscopy (STM) detects resonance by measuring vibrational amplitude during a frequency sweep. The STM tip, mounted on a piezotube, maintains constant tunneling current and gap via feedback. At resonance, increased membrane motion reduces tip-sample distance, raises current, and triggers retraction of the tip; the resulting piezotube displacement gives the amplitude. Only the vibrational amplitude, as inferred from the average tunneling current, is recorded, since the STM’s low-noise amplifier has a cutoff frequency ($\sim~{}1$ kHz) well below the membrane’s resonance frequency. The tip is optically aligned with the membrane and gradually brought into tunneling (Fig. 3b). To minimize perturbation while maintaining displacement sensitivity, it is parked 50 $\mu$m from the clamped edge, where the membrane is stiffer than at the center.

### v2

# arXiv:2504.10579v2[quant-ph] 27 Feb 2026

@@ Abstract
The Casimir effect and superconductivity are foundational quantum phenomena whose interplay is an open question in physics, with significant implications for electron physics, quantum gravity, and high-temperature superconductivity. Determining how Casimir forces behave across a superconducting transition remains elusive due to the difficulty of realizing precise alignment, cryogenic operation, and isolating small force changes from competing effects. Recent theories predict milli-Pascal jumps in Casimir pressure across the transition, motivating experiments capable of reaching well below this regime. Here, we demonstrate an on-chip superconducting nanomechanical platform that overcomes these long-standing challenges, achieving the most parallel Casimir configurations to date. Our microchip-based parallel plates reach unprecedented area-to-separation ratios, exceeding past experiments across superconducting transitions by three orders of magnitude and yielding the strongest Casimir forces generated between compliant surfaces. Scanning tunneling microscopy (STM) directly detects the resonant motion of a suspended nanoscale plate with subatomic precision in lateral positioning and displacement, enabling suppression of van der Waals, electrostatic, and thermal effects. With verified micro-Pascal pressure resolution, our platform provides a credible entry point into a new field of quantum experiments, enabling exploration of Casimir–superconductivity interactions with the stability, parallelism, and sensitivity required to access this regime of physics.
@@ I Interplay of Casimir Forces and Superconductivity
Figure 1: Superconducting Casimir Force (a) Schematic of a superconducting Casimir cavity with two plates of area AA separated by gap $d$. Inset: plates’ superconducting transition. (b) Challenges in measuring Casimir shifts across $T_{C}$ include achieving high parallelism with compliant plates, minimally invasive readout, and isolating Casimir forces from electrostatic and measurement artifacts. (c) Plate design requirements are high compliance, high-$Q$ resonance, and high $T_{C}$ for sensitive measurements across $T_{C}$.
The Casimir effect and superconductivity are two landmark quantum phenomena that display striking macroscopic consequences of quantum mechanics. The Casimir effect [1] arises from vacuum fluctuations of electromagnetic fields and produces an attractive force between closely spaced, uncharged surfaces. As shown in Figure 1a, Casimir’s original proposal consisted of two perfectly conducting parallel plates in vacuum experiencing a force that scales as $F\propto A/d^{4}$, where $A$ is plate area and $d$ the separation. Although difficult to isolate experimentally, Casimir forces were first confirmed half a century later [2] and are now known to dominate nanoscale systems, particularly in micro-electromechanical systems. Superconductivity, discovered in 1911, represents an equally dramatic phenomenon: as shown in the resistance inset of Fig. 1a, electrons above the critical temperature $T_{C}$ behave as individual particles, while below $T_{C}$ they condense into Cooper pairs that move without resistance, forming a macroscopic quantum state with coherence lengths spanning hundreds of nanometers [3].
Together, these effects raise a central open question: how does the Casimir force change when the plates themselves become superconducting? Despite the fundamental importance of both of these well-known quantum phenomena, their interplay remains experimentally inaccessible due to their weak interaction and the stringent sensor requirements needed to measure small predicted Casimir pressure changes arising from superconductivity. The sharp onset of superconductivity at $T_{C}$ provides a natural switch to toggle this interaction, enabling precise differential measurements [4, 5]. Realizing this requires overcoming three key Casimir measurement challenges, outlined in Figure 1b.
1. Extreme parallelism: maintaining a nearly constant sub-micron gap over a large plate area at cryogenic temperatures; this demands precision alignment and nanometric feedback control [6, 7], since slight tilt strongly distorts the effective Casimir interaction.
@@
2. Minimally invasive readout across $T_{C}$: measurements above and below $T_{C}$ must preserve both the superconducting state and the Casimir signal by minimizing probe-induced noise and back-action.
@@
3. Force isolation: Casimir contributions must be separated from competing effects, including thermal expansion, material changes at $T_{C}$ [8, 9, 10], electrostatics [11, 12, 13], and measurement artifacts.
Figure 1c illustrates that the movable cavity plate used to read out Casimir force changes must satisfy several requirements: (1) high compliance to respond to weak forces; (2) high-$Q$ mechanical resonance to maximize force sensitivity; and (3) high superconducting transition temperature, which shifts superconductivity-induced changes in the electromagnetic response to higher frequencies and increases overlap with the Casimir-relevant spectrum [14]. These requirements involve coupled trade-offs, making them difficult to realize simultaneously in a single device. Together with cryogenic constraints, this has left the Casimir–superconductivity regime largely unexplored experimentally, with the most advanced prior experiments remaining orders of magnitude away from the pressure resolution needed to probe predicted Casimir shifts across $T_{C}$ [15, 16, 14].
Theoretical expectations already converge on a demanding sensitivity scale. Existing calculations for broader temperature sweeps below $T_{C}$ [14], as well as our model-agnostic survey of hybrid Drude/plasma/BCS descriptions above and below $T_{C}$ (Methods), both place expected Casimir-pressure variations in the milli-Pascal range, with some scenarios smaller. This makes sub-milli-Pascal resolution a prerequisite: without it, genuine superconductivity-related Casimir changes cannot be cleanly separated from residual systematics or used to distinguish competing baseline models.
Here we present an experimental platform capable of resolving Casimir force changes in the micro-Pascal regime across a superconducting transition, a sensitivity range that has remained experimentally inaccessible. Our approach uniquely employs scanning tunneling microscopy to read out on-chip, nanofabricated parallel-plate Casimir cavities. STM provides atomic-scale control between probe and plate, with minimal invasiveness that is difficult to achieve with other readout schemes. This allows precision cancellation of all other known sources of pressure change. Monolithic on-chip fabrication enables a degree of parallelism that is practically unachievable with precision-stage alignment, allowing the most parallel plate–plate Casimir geometry realized to date.
@@
The following sections describe the new fabrication techniques for our on-chip cavities, and introduce the STM-based detection method for nanomechanics. We then outline protocols for eliminating non-Casimir contributions and report initial calibrated frequency sweeps across $T_{C}$ to probe superconductivity-induced shifts in the Casimir force. We discuss potential theoretical implications and open questions; rather than a final answer, this work establishes a genuine entry point into the micro-Pascal regime needed to systematically explore Casimir-superconductivity physics.
@@
Figure 2: Achieving extreme parallelism on-chip. (a) Colorized SEM of a microchip with suspended NbTiN (square) membranes above a NbTiN backgate; inset shows the layered structure. Membranes feature micron-scale holes for selective plasma etching of the a-Si layer. A $1~\mu\rm m$ SiO2 layer electrically isolates the Casimir cavity. (b) A central hole in the top membrane offers visual access to the gap. (c) Close-up of 190 nm gaps sustained over millimeter-scale areas, with $\sim$153 pm sag kept taut by high-stress NbTiN. (d) Comparison of area-over-gap ratios for published Casimir setups. Here, the ideal Casimir force between perfect conductors is used as a consistent metric to quantify effective parallelism. Blue squares: plate-plate ($F\propto A/d^{4}$); red circles: sphere-plate ($F\propto R/d^{3}$). Our chip platform’s parallelism surpasses prior cryogenic and room-temperature experiments, achieving the highest Casimir force between compliant surfaces to date. See Supplementary Information J for data sources.
II Ultra-Parallel Superconducting Casimir Cavities
Over the past decade, a variety of on-chip Casimir geometries have been explored [17, 18, 19, 15, 20, 21, 22, 23]. Most conventional setups use a sphere–plate geometry for easier alignment, while plate–plate configurations, though closer to Casimir’s original theory, demand much higher parallelism and stability [6, 7, 21]. To date, no superconducting Casimir experiments have reported measurable force shifts across $T_{C}$ [15, 16, 24, 21].
[…]
@@ II Ultra-Parallel Superconducting Casimir Cavities
At 190 nm, our gap is the smallest that can be sustained over millimeter-scale areas, pushing our fabrication methods to their limit (see Methods A). This distance is also large enough to avoid complex van der Waals interactions and roughness-induced noise, both of which become significant below 100 nm [25]. The membranes combine high tensile stress, a superconducting transition temperature ($T_{C}\approx 14.2$ K), and atomically-smooth surfaces, all crucial for precise Casimir measurements. Achieving both high $T_{C}$ and high tensile stress is nontrivial, but material optimization preserves mechanical performance while ensuring a $T_{C}$ close to bulk NbTiN. This overlap with Casimir-relevant modes improves the chance of measurable effects.
[…]
@@
Figure 2 illustrates the achieved high aspect-ratio parallelism. A colorized SEM (Fig. 2a) shows suspended NbTiN membranes with micron-scale release holes; the inset gives a side-view schematic. A central hole (Fig. 2b) provides visual confirmation of suspension, though omitted in measurements where suspended modes are verified by vibrometry. Close-ups (Fig. 2c) reveal 190 nm gaps maintained over millimeter spans. Fig. 2d quantifies effective parallelism, showing our area-to-separation ratio surpasses prior Casimir experiments by orders of magnitude. Beyond extreme parallelism, the high-stress NbTiN membranes function as high-$Q$ mechanical resonators, enabling sensitive detection of Casimir-induced shifts with quality factors exceeding $10^{6}$ (Supplementary Information I).
@@
Materials engineering is central to this platform. High tensile stress keeps the membranes flat, and prevents collapse while enhancing dissipation-dilution [26] for high Q, making the membranes mechanically stable enough for STM interrogation. Years of cleanroom optimization produced films with both high $T_{C}$ and resistance to processing chemicals (see Supplementary Information B). Crucially, NbTiN’s minimal oxidation allows effective in-situ annealing for tunneling, unlike aluminum which forms thick oxides. These combined properties enable both the generation of strong, well-defined Casimir forces and the atomically clean, mechanically stable surfaces required for scanning tunneling microscopy, as detailed in the next section.
III Experimental Setup and Measurement Protocol
@@
Most Casimir experiments rely on optical or capacitive readout, which interrogates millions of atoms on a plate and can disturb superconductivity through heating or electromagnetic noise. While superconducting LC resonators offer extreme sensitivity [18, 23], they only operate below $T_{C}$, cannot probe changes across the transition, and require complex fabrication. Optical approaches [15, 16] also struggle near $T_{C}$ due to local absorptive heating and poor thermal anchoring in nanoscale suspended structures. To resolve superconductivity’s influence on Casimir forces, a minimally invasive method is desirable – ideally avoiding optics, magnetic fields, and coupling only to a few atoms on the membrane at a time.
[…]
@@
Figure 3a shows the experimental setup with detailed procedure in Methods and Supplementary Information. Measurements are conducted in ultra-high vacuum ($\sim$$10^{-10}$ mbar) and cryogenic temperatures (4.45–14.45 K). A suspended NbTiN membrane vibrates above a fixed NbTiN backgate, driven by an AC+DC voltage (AC for actuation, DC to cancel electrostatics). The STM tip is optically aligned with the membrane, gradually brought into tunneling near its stiffer edge to minimize perturbation while maintaining sensitivity (Fig. 3c), and used to detect resonance by measuring tunneling current during a frequency sweep. The tip, mounted on a piezotube, maintains constant tunneling current and gap via feedback. At resonance, increased membrane motion modulates tip-sample distance, which influences the tunneling current. Since the STM’s low-noise amplifier has a cutoff frequency ($\sim$1 kHz) well below the membrane’s resonance frequency, only the average current is recorded. Due to the exponential dependence of the tunneling current on the tip-sample distance, the oscillation results in an increase of the average current, triggering a retraction of the tip. The resulting piezotube displacement follows the oscillation amplitude.